# What To Do (and Not to Do) with Causal Panel Analysis under Parallel Trends:

### v2

@@ Abstract
Two-way fixed effects (TWFE) models are ubiquitous in causal panel analysis in political science. However, recent methodological discussions challenge their validity in the presence of heterogeneous treatment effects (HTE) and violations of the parallel trends assumption (PTA). This burgeoning literature has introduced multiple estimators and diagnostics, leading to confusion among empirical researchers on two fronts: the reliability of existing results based on TWFE models and the current best practices. To address these concerns, we examined, replicated, and reanalyzed 37 articles from three leading political science journals that employed observational panel data with binary treatments. Using six newly introduced HTE-robust estimators, we find that although precision may be affected, the core conclusions derived from TWFE estimates largely remain unchanged. PTA violations and insufficient statistical power, however, continue to be significant obstacles to credible inferences. Based on these findings, we offer recommendations for improving practice in empirical research.
Keywords: two-way fixed effects, parallel trends, panel data, heterogeneous treatment effects, pretrend, difference-in-differences
@@ 1.  Introduction
in which $Y_{i,t}$ is the outcome variable for unit $i$ at time $t$; $D_{i,t}$ is the treatment variable; $X_{i,t}$ is a vector of covariates; $\alpha_{i}$, $\xi_{t}$ are unit fixed effects and time fixed effects; and $\varepsilon_{i,t}$ are idiosyncratic errors. [^2] This is with some abuse of notation. In some articles that we classify as using TWFE models, “unit” fixed effects are at the level of “groups” $g$, within which multiple units $i$ are nested (e.g. county fixed effects when $i$ indexes city), or time fixed effects are at a higher level $p$ (e.g. year fixed effects when $t$ indexes day). While it would be more general and accurate to use the notation $\alpha_{g}$ and $\xi_{p}$, with $g=i$ and $p=t$ when fixed effects are at the finest-unit or -time level, we elect not to for the sake of simplicity. The primary parameter of interest is $\delta^{TWFE}$, which researchers often interpret as the treatment effect. Moreover, researchers often equate TWFE models such as Model (1) with difference-in-differences (DID) designs and use the two terms interchangeably.
Recent methodological discussions challenge the validity of TWFE models, leaving empirical researcher in a quandary. First, they do not know whether existing findings based on TWFE models are trustworthy. Second, with the introduction of numerous new estimators and diagnostics, researchers are uncertain about which estimator is most appropriate for their research context and what diagnostics they should employ. This paper seeks to bridge this gap by reviewing new methods from the methodological literature and by reanalyzing published studies using both these new methods and the traditional TWFE models. Based on the findings, we offer recommendations for researchers.
@@
This paper thus pursues two goals. First, we explain and compare six recent proposals to amend TWFE models, including the interaction weighted (IW) estimator (Sun and Abraham, 2021), stacked DID (Baker, Larcker and Wang, 2022), CS-DID (Callaway and Karami, 2023), DID multiple (de Chaisemartin and D’Haultfœuille, 2017), PanelMatch (Imai, Kim and Wang 2023, hereafter IKW 2023), and the imputation method (Borusyak, Jaravel and Spiess 2023, hereafter BJS 2023; Liu, Wang and Xu 2022, hereafter LWX 2022). These estimators produce causally interpretable estimates under HTE and the PTA or its variants. Second, we replicate and reanalyze 37 studies published in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and Journal of Politics (JOP) from 2017 to 2022 and whose main identification strategies rely on the PTA. Our aim is to assess the consequences of using or not using HTE-robust estimators and the severity of PTA violations in political science research.
Our reanalysis shows that HTE-robust estimators largely produce point estimates in alignment with those from TWFE models. In fact, these estimators only change the sign of the original findings in a single study, suggesting that the alarm over HTE generated by the theoretical literature is potentially out of proportion. The HTE-robust estimates are, however, more often statistically insignificant than the TWFE estimates. This may be attributed to the loss of efficiency that arises when switching to HTE-robust estimators, as well as the potential overconfidence of the TWFE estimator. Moreover, we also observe signs of severe PTA violations in a large number of studies, which likely have led to spurious findings. This underlines how PTA violations persist in threatening the credibility of inferences drawn from observational panel data despite being a long- and well-known assumption.
Overall, we find that 30-40% of the articles in our sample present compelling evidence, based on current methodological literature, that the PTA is plausible and the average treatment effect on the treated (ATT) is statistically distinguishable from zero. This is not to say that the remaining studies lack credibility; often, the data at hand do not allow us to detect or dismiss violations of the PTA or to reject the null hypothesis of no effects with sufficient power when using an HTE-robust estimator.
@@
These findings have important implications for practice. First, whenever the PTA is invoked, it is essential for researchers to evaluate its plausibility using both graphical and statistical tools. Second, we advise researchers to favor HTE-robust estimators over TWFE. We argue that the potential loss of efficiency is outweighed by the risk of potential biases arising from HTE. The specific choice of estimator will depend on the research context and feasibility.
@@
In addition to providing researchers with a set of practical recommendations, this paper makes several contributions. First, we propose a typology for and provide a comprehensive comparison of various estimators for causal panel analysis with binary treatments under the PTA or its variants. We discuss the properties and assumptions of each estimator. Moreover, our reanalysis both instills confidence in existing political science research conducted with TWFE models when done properly and raises warnings about potential risks, such as the failure of the PTA and a lack of power. Our work also contributes to the ongoing conversation on replication and reproducibility in political science (e.g., Eggers et al., 2015; Lall, 2016; Hainmueller, Mummolo and Xu, 2019; Lal et al., 2023).
@@
Our work is closely related to Baker, Larcker and Wang (2022), who evaluate the credibility of a handful of studies with staggered adoption designs in the finance and accounting literature. Our work differs from theirs in two important ways: (1) we apply a wider range of estimators and diagnostic tests to a much larger number of empirical applications from a more diverse selection of settings, including cases with treatment reversal, and (2) this more comprehensive review finds that the weighting problem caused by HTE is not the primary threat to causal inference with panel data in political science research. Our work is also related to Roth et al. (2023), who synthesize the recent methodological advancements of DID in the econometrics literature. What differs is that we apply these innovations to empirical data and evaluate the robustness of existing findings.
@@
Our research has a few notable limitations. First, we do not explore methods that operate under sequential ignorability, an alternative identification framework that assumes no unobserved confounders but allows for dynamic treatment selection based on variables up to the current time period. Second, we do not address the challenge of cross-sectional interference, a phenomenon that is arguably prevalent in political and social settings. Third, although we point out the commonness of sensitivity to model specification in the literature by providing estimates when incorporating lagged dependent variables (LDVs) or unit-specific linear time trends (ULTs) in the Supplementary Material (SM), we do not further the debate regarding whether or why these estimates should be trusted over the original ones. Fourth, our analysis does not encompass studies that use continuous treatments, which is a common occurrence in political science research. Despite these limitations, which we aim to address in subsequent studies, our replication and reanalysis exercise provides valuable insights into some of the widespread and significant challenges in using TWFE models for panel data analysis.
@@ 2.  TWFE and Its Pitfalls
In this section, we review the pitfalls of TWFE models identified in the literature. In the classic two-group and two-period case, the OLS estimator (or equivalently, the least square dummy variable estimator) for the specification in Equation (1), $\hat{\delta}^{TWFE}$, is equivalent to the DID estimator, which consistently estimates the ATT under the PTA even with HTE. These results do not hold more generally in more complex settings with differential treatment adoption times (known as staggered adoption) or treatment reversal, as we will discuss below.
@@
A second body of research explores the implications of HTE with binary treatments within TWFE models (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Callaway and Sant’Anna, 2021; Borusyak, Jaravel and Spiess, 2023; Athey and Imbens, 2022). Most of this literature assumes staggered adoption, but the insights from that setting are still relevant when there are treatment reversals. In Figure 1, we present two simplified examples with staggered treatment adoption. Figure 1(a) represents outcome trajectories in line with standard TWFE assumptions, which not only include the PTA but also require that the treatment effect be immediate and unvarying across units and over time. In contrast, the right panel portrays a scenario where the PTA holds but the constant treatment effect assumption is not met. Various decompositions by the aforementioned researchers reveal that even under the PTA, when treatments begin at different times (such as in staggered adoption) and when treatment effects evolve over time, the TWFE estimand is not in general a convex combination of the ITE for observations subjected to the treatment. The basic intuition behind this theoretical result is that TWFE models use post-treatment data from units who adopt treatment earlier in the panel as controls for those who adopt the treatment later (e.g., Goodman-Bacon, 2021). HTE-robust estimators capitalize on this insight by avoiding these ‘invalid’ comparisons between two treated units. [^3] The comparisons are valid if we are willing to assume constant treatment effects but not otherwise. In the next section, we present a typology of HTE-robust estimators, along with an introduction to the estimators themselves.
@@
A third limitation of the canonical TWFE model is its presumption of no temporal and spatial interference. In most uses of TWFE models, researchers assume that there are no spatial spillovers and that treatment effects occur contemporaneously, hence no anticipation or carryover effects. [^4] No anticipation effects means that future treatments do not affect today’s potential outcomes; no carryover effects means that today’s treatment does not affect future potential outcomes. These are obviously strong assumptions that are rarely questioned or tested in practice (Imai and Kim, 2019; Athey and Imbens, 2022; Wang, 2021). Although some recent methods permit arbitrary carryover effects in staggered adoption settings (Strezhnev, 2018; Callaway and Sant’Anna, 2021), they are not distinguishable from contemporaneous effects. This limitation becomes more complex when treatment reversal is possible, as demonstrated in Figure 1. In Figure 1(b), data in the left panel are consistent with TWFE assumptions, while the right panel shows deviations from the PTA, constant treatment effect, and the absence of anticipation or carryover effects. What is concerning is that many real-world data resemble the problematic right figure rather than the ideal left one. Nevertheless, recent methods have been proposed to handle arbitrary carryover effects over a limited number of periods in more general settings (IKW 2023; LWX 2022). The challenge of addressing spatial spillover effects without strong structural assumptions still persists (Aronow, Samii and Wang, 2020; Wang, 2021), but its resolution is beyond the scope of this paper.
@@
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^6] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$) as $\delta_{l}$ ($\delta_{g,l}$) and use $l=1$ to represent the period immediately after treatment adoption. [^7] Another common practice used by some authors we reference is to denote this first post-treatment period with $l=0$. $\delta_{g,l}$ is also what some authors refer to as cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
Each of the estimators we discuss can be used to estimate $\delta_{l}$. The analogous specification for estimating DTE is a lags-and-leads specification. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods since the most recent event date that unit $i$ has been in treatment at period $t$ ($K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
where $a$ and $b$ are the number of lag and lead terms (BJS 2023). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period as suggested by Roth (2022). Conventionally, $\hat{\delta}^{TWFE}_{l}$ is interpreted as an estimate of $\delta_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\delta}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect. We refer readers to the section on “Implementation Details” in the SM for more information on Model (3), including in the case where there are treatment reversals.

### v3

@@ Abstract
Two-way fixed effects (TWFE) models are ubiquitous in causal panel analysis in political science. However, recent methodological discussions challenge their validity in the presence of heterogeneous treatment effects (HTE) and violations of the parallel trends assumption (PTA). This burgeoning literature has introduced multiple estimators and diagnostics, leading to confusion among empirical researchers on two fronts: the reliability of existing results based on TWFE models and the current best practices. To address these concerns, we examined, replicated, and reanalyzed 37 articles from three leading political science journals that employed observational panel data with binary treatments. Using six newly introduced HTE-robust estimators, along with diagnostics tests and uncertainty measures that are robust to PTA violations, we find that only a small minority of studies are highly robust. Although HTE-robust estimates tend to be broadly consistent with TWFE estimates, discrepancies in point estimates, increased measures of uncertainty, and potential PTA violations call into question many results that were already on the margins of statistical significance. We offer recommendations for improving practice in empirical research based on these findings.
Keywords: panel data, two-way fixed effects, parallel trends, heterogeneous treatment effects, pretrend, event-study plot, difference-in-differences, robust confidence set
@@ 1.  Introduction
in which $Y_{i,t}$ is the outcome variable for unit $i$ at time $t$; $D_{i,t}$ is the treatment variable; $X_{i,t}$ is a vector of covariates; $\alpha_{i}$, $\xi_{t}$ are unit fixed effects and time fixed effects; and $\varepsilon_{i,t}$ are idiosyncratic errors. [^2] This is with some abuse of notation. In some articles that we classify as using TWFE models, “unit” fixed effects are at the level of “groups” $g$, within which multiple units $i$ are nested (e.g. county fixed effects when $i$ indexes city), or time fixed effects are at a higher level $p$ (e.g. year fixed effects when $t$ indexes day). While it would be more general and accurate to use the notation $\alpha_{g}$ and $\xi_{p}$, with $g=i$ and $p=t$ when fixed effects are at the finest-unit or -time level, we elect not to for the sake of simplicity. The primary parameter of interest is $\delta^{TWFE}$, which researchers often interpret as the treatment effect. Moreover, researchers often equate TWFE models such as Model (1) with difference-in-differences (DID) designs and use the two terms interchangeably. [^3] We use the word “design” in reference to research design (King, Keohane and Verba, 2021), which differs from the typical usage in the causal inference literature that refers to treatment assignment mechanism.
Recent methodological discussions challenge the validity of TWFE models, leaving empirical researchers in a quandary. First, they do not know whether existing findings based on TWFE models are trustworthy. Second, with the introduction of numerous new estimators and diagnostics, researchers are uncertain about which estimator is most appropriate for their research context and what diagnostics they should employ. This paper seeks to bridge this gap by reviewing new diagnostic, estimation, and uncertainty quantification methods from the methodological literature and by reanalyzing published studies using both new estimators and the traditional TWFE model. Based on the findings, we offer recommendations for researchers.
@@
This paper thus pursues two goals. First, we explain and compare six recent proposals to amend TWFE models, including the interaction weighted (IW) estimator (Sun and Abraham, 2021), stacked DID (Baker, Larcker and Wang, 2022), CS-DID (Callaway and Karami, 2023), DID multiple (de Chaisemartin and D’Haultfœuille, 2017), PanelMatch (Imai, Kim and Wang 2023, hereafter IKW 2023), and the imputation method (Borusyak, Jaravel and Spiess 2023, hereafter BJS 2023; Liu, Wang and Xu 2022, hereafter LWX 2022). These estimators produce causally interpretable estimates under HTE and the PTA or its variants. Second, we replicate and reanalyze 37 studies published in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and The Journal of Politics (JOP) from 2017 to 2022 whose main identification strategies rely on the PTA. Our aim is to assess the consequences of using or not using HTE-robust estimators and of potential PTA violations in political science research.
Our reanalysis shows that, although theoretically possible, HTE-robust estimators rarely result in complete reversals of conclusions based on the TWFE estimator in the political science literature. In fact, we observe only a single study where any of these estimators change the sign of the original finding, suggesting that the alarm over HTE generated by the theoretical literature may be out of proportion. The danger is instead more subtle: HTE-robust estimators tend to have larger measures of uncertainty, which, combined with even small fluctuations in point estimates, can undermine our confidence in a result. This is especially relevant for results that originally teeter on the brink of significance.
Moreover, many studies either show clear signs of PTA violations or lack the power to dismiss the possibility that PTA violations of realistic sizes are responsible for a non-zero estimate of an average causal effect. In some studies, we observe signs of severe PTA violations, which likely have led to spurious findings. In others, allowing even mild violations of the PTA causes the original treatment effect estimates to become statistically indistinguishable from zero. This finding underscores how PTA violations persist and threaten the credibility of inferences drawn from observational panel data, despite being a longstanding and well-known assumption.
Overall, we find that about one-fourth to one-third of the articles in our sample present compelling empirical evidence to support their findings. These articles provide sufficient evidence that the PTA is plausible, and the statistically significant findings are robust under HTE and mild violations of the PTA benchmarked against observed “pretrends,” which are differences-in-means in the observed and predicted treated outcomes in the pretreatment periods. This does not imply that the remaining studies are incorrect; often, the available data do not permit detection or dismissal of PTA violations, nor do they allow us to reject the null hypothesis of no effects with sufficient power when using an HTE-robust estimator.
@@
These findings have important implications for practice. First, whenever the PTA is invoked, it is essential for researchers to evaluate its plausibility using both graphical and statistical tools, as well as to consider the consequences of its potential violations through sensitivity analyses. Importantly, showing that coefficients of the “pretrend” are not statistically significantly different from zero does not suffice to validate the PTA. This is not only because the PTA involves counterfactuals and is inherently untestable, but also a classical case of equating “absence of evidence” as “evidence of absence”—the lack of statistical significance could simply result from a lack of power. Second, we recommend that researchers favor HTE-robust estimators over TWFE. This is because the potential loss of efficiency, we believe, is justified by the reduction in risk of biases arising from HTE. We discuss the specific choice of estimator, which will depend on the research context and feasibility, in the Supplementary Materials (SM).
@@
In addition to offering researchers a set of practical recommendations, this paper makes several contributions. First, we propose a typology for and provide a comprehensive comparison of various estimators for causal panel analysis. We hope this discussion helps researchers deepen their understanding of these estimators. Ultimately, the validity of the identifying assumptions is more crucial than the choice of estimator, provided that the estimator is appropriate for the specific research context. Second, we adapt the robust confidence set (CS) approach for sensitivity analysis proposed by Rambachan and Roth (2023) to the setting of counterfactual estimators. Third, our reanalysis both instills confidence in existing political science research conducted with TWFE models when done properly and raises warnings about potential risks, such as the failure of the PTA and a lack of power. Our work also contributes to the ongoing conversation on replication and reproducibility in political science (e.g., Eggers et al., 2015; Lall, 2016; Hainmueller, Mummolo and Xu, 2019; Lal et al., 2024).
@@
Our work is closely related to Baker, Larcker and Wang (2022), who evaluate the credibility of a handful of studies with staggered adoption treatments in the finance and accounting literature. It differs from theirs in two important ways: (1) we apply a wider range of estimators and diagnostic tests to a much larger number of empirical applications from a more diverse selection of settings; his includes cases with treatment reversal, to which the majority of the studies we replicated belong. (2) This more comprehensive review suggests that the weighting problems caused by HTE may not be the primary threat to causal inference with panel data in political science research; instead, the primary concerns are PTA violations and a lack of statistical power. Our work is also related to Roth et al. (2023), Xu (2023), and Arkhangelsky and Imbens (2023), who review and synthesize the recent methodological advancements of DID in the methodological literature. What differs is that we apply these innovations to empirical data, evaluate the robustness of existing findings, and provide practical advice to researchers.
@@
Our research has a few limitations. First, we do not explore methods that operate under sequential ignorability, an alternative identification framework that assumes no unobserved confounders but allows for dynamic treatment selection based on variables up to the current time period. Second, we do not address the challenge of cross-sectional interference, a phenomenon that is arguably prevalent in political and social settings. Third, although we point out how commonly the incorporation of lagged dependent variables (LDVs) or unit-specific linear time trends (ULTs) alters results in the SM, furthering the debate regarding whether or why these estimates should be trusted over the original ones is beyond the scope of this paper. Fourth, our analysis does not encompass studies that use continuous treatments, which is a common occurrence in political science research. Despite these limitations, which we aim to address in subsequent studies, our replication and reanalysis exercise provides valuable insights into some of the widespread and significant challenges in using TWFE models for panel data analysis.
@@ 2.  TWFE and Its Pitfalls
In this section, we review the pitfalls of TWFE models identified in the literature. In the classic two-group and two-period case, the OLS estimator (or equivalently, the least square dummy variable estimator) for the specification in Equation (1), $\hat{\delta}^{TWFE}$, is equivalent to the DID estimator, which consistently estimates the average treatment effect on the treated (ATT) under the PTA even with HTE. These results do not hold more generally in more complex settings with differential treatment adoption times (known as staggered adoption) or treatment reversal, as we will discuss below.
@@
A second body of research explores the implications of HTE with binary treatments within TWFE models (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Callaway and Sant’Anna, 2021; Borusyak, Jaravel and Spiess, 2023; Athey and Imbens, 2022). Most of this literature assumes staggered adoption, but the insights from that setting are still relevant when there are treatment reversals. In Figure 1, we present two simplified examples with staggered treatment adoption. Figure 1(a) represents outcome trajectories in line with standard TWFE assumptions, which not only include the PTA but also require that the treatment effect be immediate and unvarying across units and over time. In contrast, the right panel portrays a scenario where the PTA holds, but the constant treatment effect assumption is not met. Various decompositions by the aforementioned researchers reveal that even under the PTA, when treatments begin at different times (such as in staggered adoption) and treatment effects evolve over time, the TWFE estimand is not in general a convex combination of the ITE for observations subjected to the treatment. The basic intuition behind this theoretical result is that TWFE models use post-treatment data from units who adopt treatment earlier in the panel as controls for those who adopt the treatment later (e.g., Goodman-Bacon, 2021). HTE-robust estimators capitalize on this insight by avoiding these “invalid” comparisons between two treated units in post-treatment periods. In the next section, we present a typology of HTE-robust estimators, along with an introduction to the estimators themselves.
@@
A third limitation of the canonical TWFE model is its presumption of no temporal and spatial interference. In most uses of TWFE models, researchers assume that there are no spatial spillovers and that treatment effects occur contemporaneously, hence no anticipation or carryover effects. [^4] No anticipation effects means that future treatments do not affect today’s potential outcomes; no carryover effects means that today’s treatment does not affect future potential outcomes. These are obviously strong assumptions that are rarely questioned or tested in practice (Imai and Kim, 2019; Athey and Imbens, 2022; Wang, 2021). Although some recent methods permit arbitrary carryover effects in staggered adoption settings (Strezhnev, 2018; Callaway and Sant’Anna, 2021), they are not distinguishable from contemporaneous effects. [^5] See Section A.4. of the SM in Liu, Wang and Xu (2022) for more details. This limitation becomes more complex when treatment reversal is possible, as demonstrated in Figure 1. In Figure 1(b), data in the left panel are consistent with TWFE assumptions, while the right panel shows deviations from the PTA, constant treatment effect, and the absence of anticipation or carryover effects. What is concerning is that many real-world data resemble the problematic right figure rather than the ideal left one. Nevertheless, recent methods have been proposed to handle arbitrary carryover effects over a limited number of periods in more general settings (IKW 2023; LWX 2022). The challenge of addressing spatial spillover effects without strong structural assumptions still persists (Aronow, Samii and Wang, 2020; Wang, 2021), but its resolution is beyond the scope of this paper.
@@
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^7] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$) as $\delta_{l}$ ($\delta_{g,l}$) and use $l=1$ to represent the period immediately after treatment adoption. [^8] Some authors we reference to denote this first post-treatment period with $l=0$. $\delta_{g,l}$ is also what some authors refer to as cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
Each of the estimators we discuss can be used to estimate $\delta_{l}$. The specification analogous to TWFE for estimating DTE is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}/leq0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
where $a$ and $b$ are the number of lag and lead terms (BJS 2023). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\delta}^{TWFE}_{l}$ is interpreted as an estimate of $\delta_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\delta}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect. We refer readers to Section A.4 in the SM for more information on Model (3), including in the case where there are treatment reversals.